The sum of three consecutive odd intergers is 40 more than the smallest. What are the integers?
Answer by mananth(16946)

let the odd integers be x,x+2,x+4
x+x+2+x+4 = x+40
3x+6 =x+40
2x=34
x= 17
..
The numbers are 17,19,21
..
m.ananth@hotmail.ca